OpenStudy (anonymous):

Write the standard equation for the circle. center (10, –6), r = 6 A. (x + 10)^2 + (y – 6)^2 = 6 B. (x + 6)^2 + (y – 10)^2 = 36 C. (x – 10)^2 + (y + 6)^2 = 36 D. (x – 10)^2 + (y + 6)^2 = 6

OpenStudy (anonymous):

equation for circle (x - h)^2 + (y-k)^2 = r^2 where h is x-coordinate of center, and k is y-coordinate of center. Just plug everything in now

OpenStudy (anonymous):

okay , one second

OpenStudy (anonymous):

okay so it would now look like , (x-10^2+(y-6)^2=r^2?

OpenStudy (anonymous):

so close y+6 not y-6 (because y-(-6) is y+6) and plug in 6 for r

OpenStudy (anonymous):

okay , so now it looks like , x-10^2+(y+6)^2=6^2?

OpenStudy (anonymous):

yeah but remember your parentheses and you can write 36 instead of 6^2 at the end: (x–10)^2 + (y+6)^2 = 36

OpenStudy (anonymous):

right, :) ,which would mean answer choice C , thanks so much! , but may i ask how you know to plug in h for x and k for y?

OpenStudy (anonymous):

np as for h and k - the standard equation for a circle is (x - h)^2 + (y-k)^2 = r^2 as I've said before. If you know the formula, you know what h and k are. (I'm surprised you were given the problem without the formula)

OpenStudy (anonymous):

oh , sorry forgot you stated that before lol , but yeah that's definitely my major problem is know the formulas or which one to use because they never tell us :/

OpenStudy (anonymous):

*knowing

OpenStudy (anonymous):

haha well now you know